# Gravitational waves from two scalar fields unifying the dark sector with inflation

### v1

@@ I Introduction
Since the striking discovery that the Universe is presently undergoing a second period of accelerated expansion [6, 7], driven by a still unknown form of energy, dubbed dark energy, several attempts have been made to unify within a single theoretical framework both inflation and this energy. If dark matter—whose nature remains currently elusive—is also included, then one can imagine that both the dark sector, composed of dark energy and dark matter, plus inflation, may be framed in terms of a triple unification cosmological scheme [8, 9, 10, 11, 12, 13].
@@ II Two-scalar-field cosmological model
where $a(t)$ is the scale factor, $t$ is the cosmic time and $d\Sigma^{2}$ is the metric of the three-dimensional Euclidean space. In this way, the equations of motion during the first stage of evolution (inflation and smooth transition to a radiation-dominated era) are
@@
During the second stage of evolution (radiation-, matter-, and dark energy-dominated eras), the corresponding evolution equations are
@@
where $C$ is a dimensional constant whose value is fixed by current cosmological measurements, and further we have introduced ordinary baryonic matter, described as a perfect fluid with pressure $p_{\texttt{BM}}=0$ and energy density $\rho_{\texttt{BM}}=\rho_{\texttt{BM}0}\,e^{-3u}$.
@@ III Formalism of continuous Bogoliubov coefficients
For $k^{2}\lesssim a^{\prime\prime}/a$, the system of ordinary differential equations (18) describes the production of gravitational waves with angular frequency $\omega=k/a$, while for $k^{2}\gg a^{\prime\prime}/a$, it describes a harmonic oscillator, implying that no gravitational waves are produced. This circumstance will be utilized in our numerical calculations (see Sect. IV) to reduce the computation time of gravitational wave spectra substantially.
@@ IV Gravitational-wave spectrum
Using the formalism of continuous Bogoliubov coefficients, we now calculate the full energy spectrum of the gravitational waves generated within the two-scalar-field cosmological model described in Sect. II.
To this end, we numerically solve the system of equations (20) for a specific value of $\omega_{0}$, with initial conditions corresponding to the absence of gravitational waves ($X_{i}=Y_{i}=1$). In this system of equations, the values of $\dot{a}/a$ and $\ddot{a}/a$ as functions of $u$ are known from the beginning of inflation to the present moment (see Sect. II). In this way, we obtain $X$ and $Y$ as functions of $u$, which, in turn, allows us to obtain $|\beta_{k}|^{2}$ as a function of $u$, and, in particular, its value at the present moment, $(|\beta_{k}|^{2})_{0}$. Finally, using Eq. (19), we calculate $\Omega_{\texttt{GW}}$ for the value of $\omega_{0}$ under consideration. Repeating this procedure for multiple values of $\omega_{0}$ between $\omega_{\texttt{min}}$ and $\omega_{\texttt{max}}$, we obtain the full energy spectrum of gravitational waves.
@@
In Fig. 3, we present the full gravitational-wave energy spectrum for the base scenario of the two-scalar-field cosmological model, whose parameter values and initial conditions were specified at the end of Sect. II. The gravitational wave density parameter $\Omega_{\texttt{GW}}$ is presented as a function of the frequency $f$, related to the angular frequency $\omega_{0}$ by the relation $\omega_{0}=2\pi f$. The sensitivity curves of planned ground- and space-based gravitational-wave detectors—Laser Interferometer Space Antenna (LISA), Big Bang Observer (BBO), Cosmic Explorer (CE), Einstein Telescope (ET), Square Kilometre Array (SKA), International Pulsar Timing Array (IPTA), and Deci-Hertz Interferometer Gravitational Wave Observatory (DECIGO)—are superimposed on the spectrum [^7] For further details regarding the sensitivity curves and characterization of the detectors, see Refs. [45, 46].
@@ V Final outlooks
Hence, we presented the energy density spectrum of gravitational waves, $\Omega_{\mathrm{GW}}(f)$, as a function of frequency and systematically compared it with the projected sensitivities of future ground- and space-based detectors, including LISA, BBO, Cosmic Explorer, Einstein Telescope, SKA, IPTA, and DECIGO.
@@ Acknowledgements.
OL acknowledges support by the Fondazione ICSC, Spoke 3 Astrophysics and Cosmos Observations. National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) Project ID $CN00000013$ “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19)” - Next Generation EU (NGEU). TM acknowledges hospitality to the University of Algarve during the time in which this paper has been finalized. TM is also grateful to Youri Carloni for fruitful discussions. PS acknowledges support from Fundação para a Ciência e a Tecnologia (Portugal) through the research grant UID/4106/2025 (CIDMA-Center for Research and Development in Mathematics and Applications) and the program contract UID/PRR/4106/2025.

### v2

@@ I Introduction
Since the striking discovery that the Universe is presently undergoing a second period of accelerated expansion [6, 7], driven by a still unknown form of energy, dubbed dark energy, several attempts have been made to unify within a single theoretical framework both inflation and this energy. If dark matter—whose nature remains currently elusive—is also included, then one can imagine that both the dark sector, composed of dark energy and dark matter, plus inflation, may be framed in terms of a triple unification cosmological scheme [8, 9, 10, 11, 12, 13, 14, 15].
@@ II Two-scalar-field cosmological model
where $a(t)$ is the scale factor, $t$ is the cosmic time, and $d\Sigma^{2}$ is the metric of the three-dimensional Euclidean space. In this way, the equations of motion during the first stage of evolution (inflation and smooth transition to a radiation-dominated era) are
@@
To characterize the intensity of the energy transfer from the inflaton field to the radiation bath during the inflationary period, it is common to introduce the dissipation ratio $Q$, defined as $Q\equiv\Gamma/(3H)$. Depending on the value of $Q$, two regimes of warm inflation can be defined: the weak dissipative regime ($Q<1$) and the strong dissipative regime ($Q>1$). The two-scalar-field cosmological model under consideration can accommodate both regimes, as shown in Ref. [17]. In this paper, we choose a set of values for the model’s parameters and initial conditions (see below), yielding a strong dissipative regime.
At the end of the inflationary period, the dissipation coefficients $\Gamma_{\xi}$ and $\Gamma_{\phi}$ are exponentially suppressed and, soon afterward, become negligible. This marks the end of the first stage of evolution. During the second stage, in the absence of dissipation, radiation decouples from the scalar field $\xi$, which begins to oscillate rapidly around the minimum of its quadratic potential [^2] The oscillatory regime is ensured by the condition $M_{\xi}(\phi)\gg H$, where $M_{\xi}^{2}(\phi)\equiv m^{2}e^{-\beta\kappa\phi}$ is the effective mass of the inflaton field., behaving like a nonrelativistic dark-matter fluid with equation of state $p_{\xi}=0$ and energy density given by [17]
$\rho_{\xi}=Ce^{-3u}e^{\frac{(\alpha-\beta)}{a}\phi},$ (8)
where $C$ is a constant whose value is fixed by current cosmological measurements. The previous expression reveals that in the presence of a direct coupling between the scalar fields $\xi$ and $\phi$, dark matter does not evolve like usual baryonic matter ($e^{-3u}$, or $a^{-3}$ in terms of the scale factor). There is now an explicit dependence on the dark energy field $\phi$, requiring dark matter and baryonic matter to be treated separately.
Accordingly, during the second stage of evolution (radiation-, matter-, and dark energy-dominated eras), the evolution equations are
@@
where ordinary baryonic matter is described as a perfect fluid with pressure $p_{\texttt{BM}}=0$ and energy density $\rho_{\texttt{BM}}=\rho_{\texttt{BM}0}\,e^{-3u}$.
@@
The above choice of parameters and initial conditions corresponds to a strong dissipative regime ($Q>1$). In Fig. 1, the evolution of $Q$ is shown for the entire inflationary period, as well as for the smooth transition from inflation to the radiation-dominated era, while Fig. 2 shows the evolution of the ratios $T/H$, $\rho_{\texttt{R}}/\rho_{\xi}$, and $\rho_{\texttt{R}}/\rho_{\phi}$ during inflation.
Figure 1: Evolution of the dissipation ratio $Q$ during the inflationary period, which extends from $u\approx-135.3$ to $u\approx-65.3$ (7070 e-folds of expansion), and the transition from inflation to the radiation-dominated era, which extends from $u\approx-65.3$ to $u\approx-64.3$. Throughout the entire inflationary period, a strong dissipative regime is maintained. At the end of this period, the dissipation coefficients $\Gamma_{\xi}=\Gamma_{\phi}$ are exponentially suppressed, resulting in a sharp decrease of $Q$.
Figure 2: Evolution of the ratios $T/H$, $\rho_{\texttt{R}}/\rho_{\xi}$, and $\rho_{\texttt{R}}/\rho_{\phi}$ during the inflationary period. The end of inflation occurs when $\rho_{\texttt{R}}/\rho_{\xi}=1$. From then on, radiation dominates the Universe’s evolution.
Finally, it is noteworthy to focus on the cosmological perturbations induced by such a model. During the inflationary epoch, scalar and tensor perturbations are expected to be relevant and detectable. The latter are studied in this article. As for the former, they will be analyzed in detail in a future publication; here, we will limit our comments to some general observations. Scalar perturbations are fundamental to validate cosmological models by comparing their predictions with the observational constraints provided by the Planck mission [27]. The two-scalar-field model under consideration presents a complex scenario in this regard; indeed, considering warm inflation within a multi-field approach leads to an unknown picture for describing scalar perturbations (as far as we know). In warm inflation, the source of density fluctuations is the thermal fluctuations in the radiation bath, and this is a physically different scenario with respect to cold inflation, where the source is the quantum fluctuations of the inflaton field [23, 20]. On the other hand, multi-field approaches provide a more complex framework, where the interaction among the fields generally generates scalar perturbations that are superpositions of several contributions, such as entropic and adiabatic ones [28]. Since both warm inflation and multi-field inflation have proven to be suitable frameworks, as shown in Refs. [29, 30, 31, 32, 33, 34], it is expected that the two-scalar-field cosmological model can be successfully constrained as well. In particular, for the choice of parameters in the base scenario, since $\rho_{\phi}\ll\rho_{\xi}$ (see Fig. 2), standard single-field warm inflation results should provide a good approximation. Thus, as shown in Refs. [30, 22], considering a strong dissipation regime should lead to observational consistency, with a reduced tensor-to-scalar ratio and a more blue-tilted spectrum as compared to cold inflation. Nevertheless, to properly account for the predictions of the two-scalar-field cosmological model and to rigorously test its consistency, the generalization of the aforementioned approaches should be considered. However, this is not straightforward and, as mentioned above, is left for future work.
@@ III Formalism of continuous Bogoliubov coefficients
For $k^{2}\lesssim a^{\prime\prime}/a$, the system of ordinary differential equations (19) describes the production of gravitational waves with angular frequency $\omega=k/a$, while for $k^{2}\gg a^{\prime\prime}/a$, it describes a harmonic oscillator, implying that no gravitational waves are produced. This circumstance will be utilized in our numerical calculations (see Sec. IV) to reduce the computation time of gravitational wave spectra substantially.
@@ IV Gravitational-wave spectrum
Using the formalism of continuous Bogoliubov coefficients, we now calculate the full energy spectrum of the gravitational waves generated within the two-scalar-field cosmological model described in Sec. II.
To this end, we numerically solve the system of equations (21) for a specific value of $\omega_{0}$, with initial conditions corresponding to the absence of gravitational waves ($X_{i}=Y_{i}=1$). In this system of equations, the values of $\dot{a}/a$ and $\ddot{a}/a$ as functions of $u$ are known from the beginning of inflation to the present moment (see Sec. II). In this way, we obtain $X$ and $Y$ as functions of $u$, which, in turn, allows us to obtain $|\beta_{k}|^{2}$ as a function of $u$, and, in particular, its value at the present moment, $(|\beta_{k}|^{2})_{0}$. Finally, using Eq. (20), we calculate $\Omega_{\texttt{GW}}$ for the value of $\omega_{0}$ under consideration. Repeating this procedure for multiple values of $\omega_{0}$ between $\omega_{\texttt{min}}$ and $\omega_{\texttt{max}}$, we obtain the full energy spectrum of gravitational waves.
@@
In Fig. 5, we present the full gravitational-wave energy spectrum for the base scenario of the two-scalar-field cosmological model, whose parameter values and initial conditions were specified at the end of Sec. II. The gravitational wave density parameter $\Omega_{\texttt{GW}}$ is presented as a function of the frequency $f$, related to the angular frequency $\omega_{0}$ by the relation $\omega_{0}=2\pi f$. The sensitivity curves of planned ground- and space-based gravitational-wave detectors—Laser Interferometer Space Antenna (LISA), Big Bang Observer (BBO), Cosmic Explorer (CE), Einstein Telescope (ET), Square Kilometre Array (SKA), International Pulsar Timing Array (IPTA), and Deci-Hertz Interferometer Gravitational Wave Observatory (DECIGO)—are superimposed on the spectrum [^8] For further details regarding the sensitivity curves and characterization of the detectors, see Refs. [55, 56].
@@ V Final outlooks
Hence, we presented the energy density spectrum of gravitational waves, $\Omega_{\mathrm{GW}}(f)$, as a function of frequency and systematically compared it with the projected sensitivities of future ground- and space-based detectors, including LISA, BBO, CE, ET, SKA, IPTA, and DECIGO.
@@ Acknowledgements.
OL acknowledges support by the Fondazione ICSC, Spoke 3 Astrophysics and Cosmos Observations. National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) Project ID $CN00000013$ “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19)” - Next Generation EU (NGEU). TM acknowledges hospitality to the University of Algarve during the time in which this paper has been finalized. TM is also grateful to Youri Carloni for fruitful discussions. PS acknowledges support from CIDMA-Center for Research and Development in Mathematics and Applications under the Portuguese Foundation for Science and Technology Multi-Annual Financing Program for R&D Units, grants UID/04106/2025 (https://doi.org/10.54499/UID/04106/2025) and UID/PRR/04106/2025 (https://doi.org/10.54499/UID/PRR/04106/2025)